OpenStudy (anonymous):

The sum of the first n terms of an AP, 2,6,10, is 882. Find the value of n

OpenStudy (anonymous):

\[ a_i= 4(i-1) +2 \text { for } \quad i=1,2,3,\cdots \]

OpenStudy (anonymous):

\[ \sum_{i=1}^n 4(i-1) + 2 = 2n + 4( 1+2+\cdots n-1)= 288\\ 2n + 4 n(n-1)/2 = 288 2 n^2 =288\\ n^2=144\\ n=12 \]

OpenStudy (anonymous):

\[ \sum_ {i = 1}^n 4 (i - 1) + 2 = 2 n + 4 (1 + 2 + \cdots n - 1) = 288 \\ \]

OpenStudy (anonymous):

\[2 n + 4 n (n - 1)/2 = 288\implies 2 n^2 = 288\implies n^2 = 144\implies n = 12\] \]